Assume we have a fixed supply of a depletable resource to allocate between two periods. Also...

  1. Assume we have a fixed supply of a depletable resource to allocate between two periods.

Also assume that the demand function is the same in each of the two periods,

given by the formula P = 4 – 0.2q, and the marginal cost of supplying the resource is constant at $1 per unit. Now, assume the available supply of the resource q is 10 units, and the discount rate is 10%.

What is the dynamically efficient allocation of this resource? Show your work.
Answer #1

Let us check the quantity demanded at P=MC=1

P=4-0.2q

1=4-0.2q

3=0.2q

q=15

Available quantity is less than the quantity demanded, Dynamic efficient allocation is required.

Marginal Net benefit in perid 1=MNB1=P-MC=4-0.2q1-1=3-0.2q1

Marginal Net benefit in period 2=MNB2=(P-MC)/(1+r)=(3-0.2q2)/1.1

Set MNB1=MNB2

3-0.2q1=(3-0.2q2)/1.1

3.3-0.22q1=3-0.2q2

0.3=0.22q1-0.2q2 (equation 1)

We know that

q1+q2=10

So, q2=10-q1

Set q2=10-q1 in equation 1

0.3=0.22q1-0.2*(10-q1)

0.3=0.22q1-2+0.2q1

0.42q1=2.3

q1=5.476 (allocation in period 1)

q2=10-5.476=4.524 (allocation in period 2)